x^(2)+x+sqrt(x^(2)+x)-2=0
Answer by user_dude2008(1861)

x^2+x+sqrt(x^2+x)-2=0


sqrt(x^2+x)=-x^2-x+2


x^2+x=(-x^2-x+2)^2


x^2+x=x^4+2x^3-3x^2-4x+4


0=x^4+2x^3-3x^2-4x+4-x^2-x


0=x^4+2x^3-4x^2-5x+4


Answer:
x=-1/2+(1/2)*sqrt(5) or x=-1/2-(1/2)*sqrt(5)